OpenStudy (anonymous):

Can someone please help me!!! Simplify the radical expression Square root of 56x^2 1.) 28x 2.) 2x square root of 14 3.) 2x square root of 7 4.) 2 square root of 14x^2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it clear that \(\sqrt{x^2}=x\)?

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

No not really I don't understand

OpenStudy (anonymous):

Wait yes.

OpenStudy (misty1212):

do you know what the square root means?

OpenStudy (misty1212):

whew

OpenStudy (misty1212):

so the \(x\) comes outside of the radical now lets deal with \(\sqrt{56}\)

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Do we have to find the factors of it or something like that

OpenStudy (misty1212):

\[56=4\times 14\] so \[\sqrt{56}=\sqrt{4\times 14}=\sqrt4\sqrt{14}=2\sqrt{14}\]

OpenStudy (misty1212):

yeah find the perfect square that is a factor of 56 in this case it is 4

OpenStudy (misty1212):

final answer \[2x\sqrt{14}\]

OpenStudy (anonymous):

Thank you so much do you mind helping with more?

OpenStudy (misty1212):

sure \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Should I close this one or just keep going?